Grade 12th passPhysical Chemistry

100 ml of a water sample contains 40mg of Ca+2 ions . What is the hardness of water in ppm

MD hussain
7 Years agoGrade 12th pass

1 Answer

Samyak Jain
7 Years ago
Hardness of water in ppm = (mass of CaCO3 / mass of water) x 106 if hardness is caused due to Ca2+
Given mass of Ca2+ ions = 40 mg = 40 x 10–3 g
Number of moles of Ca2+ ions = 40 x 10–3 g / 40 g mol–1 = 10–3 mol
Molar mass of CaCO3 = 100 g mol–1
\therefore mass of 10–3 mol CaCO3 = 10–3 mol . 100 g mol–1 = 10–1 g
Density of water is 1 g/ml.
Thus, mass of 100 ml water is 100 g.
\therefore hardness of water in ppm = (10–1 / 100) x 106 = 103.